What is the volume of a aluminum ball at 10 degree celsius if its volume at 160 degree celsius is 300m³


1
Expert's answer
2021-03-02T18:03:35-0500
"\dfrac{L-L_0}{L_0}=\alpha_V\Delta T,""L=L_0(1+\alpha_V\Delta T),""L=L_0(1+3\alpha_L\Delta T),""L=300\ m^3\cdot(1+3\cdot23\cdot10^{-6}\ ^{\circ}C^{-1}\cdot(10^{\circ}C-160^{\circ}C))=297\ m^3."